# Alpha-element abundance patterns in star-forming regions of the local Universe

### v1

@@ 4.3 Argon
In Fig. 5, we plot log(Ar2+/O2+) (top) and log((Ar2++Ar3+)/O2+) (bottom) as a function of O2+/O, for H ii regions (blue circles) and SFGs (black squares) of the DESIRED-E sample. The fraction of objects with Ar2+ and Ar3+ abundances is 33.4% in SFGs, which is higher than in H ii regions (12.5%). This is because SFGs tend to have higher O2+/O ratios than H ii regions in galaxies (this is illustrated in Fig. 15). In Fig. 5, the coloured curved lines represent the behaviour of the different ICF(Ar) schemes considered: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colors corresponding to the three metallicity ranges) and Pérez-Montero et al. (2007, green dashed line). The horizontal black dashed line shows the value of the solar log(Ar/O) ratio –along as its uncertainty represented by the grey band– recommended by Asplund et al. (2021). The top panel of Fig. 5 shows that the curves of the different ICF(Ar) schemes reproduce fairly well the behaviour of the observational points except the one by Amayo et al. (2021), that does not reproduce the position of SFGs with O2+/O $\geq$ 0.9. In the case of the bottom panel of Fig. 5, we can see that all the curves reproduce quite well the distribution of the observational points. This is because the objects showing [Ar iv] lines in their spectra have necessarily a high ionisation degree and the contribution of unseen Ar+ should be very small. Moreover, the bottom right panel of Fig. 14 shows that the Ar abundance obtained with the three ICF(Ar) for the same object differs less than  0.04 dex, except for those with O2+/O $\geq$ 0.95.
@@ 5.1 Ne/O
According to the results analysed in Sec. 4.1 and figs. 2 and 3, the ICF(Ne) by Izotov et al. (2006) seems to better reproduce the behaviour of log(Ne/O) for H ii regions and SFGs. It shows a log(Ne/O) vs. O2+/O ratio relation flatter than the other ICF(Ne) schemes, the lowest standard deviation and a mean and median values of log(Ne/O) of --0.67 $\pm$ 0.12 and --0.66, respectively, both consistent with the solar ratio of --0.63 $\pm$ 0.06 (Asplund et al., 2021). In Sec. 4.1 we pointed out that, at values of O2+/O $<$ 0.5, an important fraction of the objects show very low log(Ne/O) values (see Fig. 3) and that almost all of them are classified as H ii regions. In Fig. 2 it is evident that the position of those points cannot be correctly reproduced by any of the ICF(Ne) schemes considered. The strong increased scatter of the Ne2+/O2+ ratio at lower O2+/O was formerly reported by Kennicutt et al. (2003). Croxall et al. (2016), in their study of H ii regions in the spiral galaxy NGC 5457 (M101), confirm a similar scatter, claiming for the existence of a large population of H ii regions with significant offset to low values of log(Ne/O).
@@
Nucleosynthesis models seem to agree in predicting a basically lockstep evolution of O and Ne (e.g. Iwamoto et al., 1999; Johnson, 2019; Kobayashi et al., 2020). In Fig. 8 we also include a curve (green line) representing the evolution of Ne and O abundances of the ISM predicted by the chemical evolution model (CEM) of the Milky Way obtained by Medina-Amayo (2023) and Amayo & Carigi (in preparation). This CEM was developed for the solar neighborhood of the Galactic disc (7--9 kpc from the Galactic centre) within a two-infall gas formation scenario, similar to that of Spitoni et al. (2021). During the first 4 Gyr, the thick disc formed from a short accretion of primordial gas. Then, the thin disc began forming from a prolonged second accretion (the loop in the curve), which has continued to the present. The star formation rate and initial mass function (IMF) are based on the formulations of Kennicutt (1998) and Kroupa (2002), respectively. The model uses stellar yields for low-and intermediate-mass stars by Ventura’s group (e.g. Ventura et al., 2022, and references therein), for SNe Ia by Leung & Nomoto (2018), and for massive stars by Nomoto et al. (2013). An upper mass limit of the IMF at 37 M was inferred to match the present-day O abundance of H ii regions obtained from CELs at $\sim$8 kpc (Méndez-Delgado et al., 2022a). In Fig. 16 we show the cumulative fraction of 16O, 20Ne, 32S, 36Ar –the most abundant stable isotopes of each element, that represent between 85 to 95% of the total amount of them– produced by different types of stars as a function of 12+log(O/H) predicted by the CEM. We distinguish the fractions produced by low- and intermediate-mass stars in the asymptotic giant branch phase (AGBs, cyan line) or type Ia supernovae (SNe Ia, dark red line) and massive stars in core collapse supernovae (CCSNe, magenta line). It is important to remember that the CEM used is built to reproduce the abundance patterns observed in the Milky Way. Objects in other galaxies with a different star-formation history may be located in somewhat different places in the abundance ratios diagrams. For example, the infall of unprocessed gas into a star-forming region in a galaxy may decrease the metallicity while keeping the abundance ratio of two alpha-elements constant.
@@
Figure 8: log(Ne/O) as a function of 12+log(Ne/H) for SFGs of the DESIRED-E sample, which Ne/O ratios have been calculated using the ICF(Ne) scheme by Izotov et al. (2006). The red continuous line represents a linear fit to the data. The green continuous line shows the time evolution of Ne and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023) and Amayo & Carigi (in preparation). The dotted red line represents the mean value of log(Ne/O). The black dashed lines and the grey bands show the solar 12+log(Ne/H) and log(Ne/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
@@
As we pointed out above, in Fig. 7, the behaviour of H ii regions is very different to that shown by SFGs. In the log(Ne/O) vs. 12+log(O/H) diagram, the observational points classified as H ii regions are distributed in a quite defined range of metallicity but with a much larger dispersion of log(Ne/O). We can distinguish that although most of the objects are concentrated at log(Ne/O) values close to the solar one, there seems to be a secondary concentration at 12+log(O/H) $\sim$ 8.4 and log(Ne/O) between --1.0 and --0.7. In fact, in Table 1 the mean log(Ne/O) of the H ii regions is 0.07–0.08 dex lower than the mean of the SFGs or the solar value. Croxall et al. (2016) and Berg et al. (2020) find a similar log(Ne/O) vs. 12+log(O/H) distribution, although this was expected given that the observational data used in those studies are included in the DESIRED-E sample. The log(Ne/O) vs. O2+/O diagram of the bottom row of Fig. 7 shows that the dispersion towards lower values of log(Ne/O) is clearly a function of O2+/O, indicating that the ICF(Ne) is not able to correct appropriately for the presence of Ne+ in H ii regions data points at O2+/O $\leq$ 0.75, which are the vast majority of the objects of this type in the DESIRED-E sample. Kennicutt et al. (2003) suggested that this behaviour is due to the high sensitivity of Ne ionisation models to the input stellar atmosphere fluxes, which in turn are very much dependent on the treatment of opacity and stellar winds. Considering the difficulties of obtaining consistent Ne/O values from Galactic and extragalactic H ii regions, we have chosen not to take them into account for the study of the behaviour of this ratio as a function of metallicity and in the data shown in Table 2.
@@
Figure 9: log(Ne/O) as a function of 12+log(O/H). Black squares and blue circles represent SFGs and H ii regions of the DESIRED-E sample, which Ne/O ratios have been calculated using the ICF(Ne) scheme by Izotov et al. (2006). The green continuous line shows the time evolution of Ne and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023) and Amayo & Carigi (in preparation). Yellow and orange stars represent the abundances obtained from quantitative spectroscopic analysis of Galactic B stars by Nieva & Przybilla (2012) and Weßmayer et al. (2022), respectively. Magenta stars represent the abundances of Galactic O stars determined by Aschenbrenner et al. (2023). The black dashed lines and the grey bands show the solar 12+log(O/H) and log(Ne/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
@@
In Fig. 9 we represent the log(Ne/O) vs. 12+log(O/H) distribution of H ii regions and SFGs of the DESIRED-E sample, the curve defined by the CEM of Medina-Amayo (2023) and Amayo & Carigi (in preparation) and abundances of O and B-type Galactic stars determined from quantitative spectroscopy by Nieva & Przybilla (2012), Weßmayer et al. (2022) and Aschenbrenner et al. (2023). The atmosphere of O and B-type stars represents the present-day chemical composition of the ISM and, in principle, should provide abundance values similar to the ionised gas associated to star-forming regions. In Fig. 9, we can see that the stellar O abundances are close to solar or oversolar –as expected because most of the stars are located in the solar neighborhood– but quite separated from the locus of the bulk of the H ii regions. A systematic difference between nebular and stellar abundances of objects located in the same star-forming regions is a well-known fact. Studies comparing the O abundances of Galactic H ii regions with that of their associated O and B-type stars (e.g. Simón-Díaz et al., 2006; Simón-Díaz & Stasińska, 2011; García-Rojas et al., 2014)) find stellar abundances $\sim$0.2 dex higher than nebular ones when they are calculated using CELs. However, such difference disappears when using faint recombination lines instead of CELs for deriving the nebular O abundance, this offset is the so-called abundance discrepancy problem. It may be related to the presence of temperature fluctuations inside the H ii regions (Peimbert, 1967; García-Rojas et al., 2007; Méndez-Delgado et al., 2023b). On the other hand, Fig. 9 indicates that the O and B-type stars also show a quite high dispersion in their log(Ne/O) ratio. The mean and standard deviation of log(Ne/O) of the stars included in the figure is --0.56 $\pm$ 0.17, a fairly supersolar value. Although temperature fluctuations can alter the abundances calculated with CELs, the abundance ratios –as Ne/O, S/O or Ar/O– obtained with this type of lines remain practically unchanged, so the observed log(Ne/O) offset between stellar and nebular objects is not expected to be related to this phenomenon (e.g. Esteban et al., 2020; Méndez-Delgado et al., 2024).
@@ 5.2 S/O
As in the case of the Ne/O ratio discussed in Sect. 5.1, the ICF(S) by Izotov et al. (2006) is the one that provides a more featureless distribution between log(S/O) vs. 12+log(O/H) and vs. O2+/O (Fig. 3), a better fit to the (S++S2+)/O vs. O2+/O distribution in the range of ionisation degrees in common with the others ICFs (Fig. 2) and a mean log(S/O) closer to the solar value of --1.57 $\pm$ 0.05 (Asplund et al., 2021). We will adopt the results obtained using this ICF(S) for the discussion we will develop in this section. The only drawback of the ICF(S) scheme by Izotov et al. (2006) is that its application is limited to objects with O2+/O $\geq$ 0.2. In the case of the DESIRED-E sample this has a very limited impact. While no SFGs have to be removed applying this limit, the fraction of H ii regions that are finally excluded is only 13.3%. Considering the rough relation between ionisation degree and metallicity shown in Fig. 15, this implies that the removed points correspond mostly to high metallicity objects but –as we can see in the panels of the left column of Fig. 4– that area is not significantly depopulated when we compare the results using the different ICF(S) schemes.
@@
Figure 11: log(S/O) as a function of 12+log(O/H) (top) and as a function of 12+log(S/H) (bottom). Black squares and blue circles represent SFGs and H ii regions of the DESIRED-E sample, which S/O ratios have been calculated using the ICF(S) scheme by Izotov et al. (2006). The green continuous line shows the time evolution of S and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023) and Amayo & Carigi (in preparation). Yellow stars represent the abundances obtained from quantitative spectroscopic analysis of Galactic B stars by Weßmayer et al. (2022). Orange stars represent the abundances of Galactic classical Cepheids determined by da Silva et al. (2023). Orange diamonds indicate the mean log(S/O) values considering bins in 12+log(S/H). The red continuous lines represent linear fits to the data represented in each panel. The black dashed lines and the grey bands show the solar 12+log(O/H), 12+log(S/H) and log(S/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
@@
The behaviour of the data points of the DESIRED-E sample shown in Figs. 10 and 11 contrasts with the results obtained by some previous studies (Dors et al., 2016; Díaz & Zamora, 2022; Brazzini et al., 2024). In particular, Díaz & Zamora (2022) find that S/O vs. O/H and vs. S/H relations are quite different for SFGs and H ii regions. Díaz & Zamora (2022) report that SFGs tend to show a S/O ratio lower than solar and that it increases as the S/H increases, but remains constant with O/H, results that are in agreement with ours. However, they find that their H ii regions show S/O ratios larger than solar and a clear tendency for lower S/O ratios as O or S abundances increase, a behaviour different to what we observe. In our case, for H ii regions, since log(S/O) remains basically constant with 12+log(O/H), it shows a clear increase as 12+log(S/H) increases (see Table 2). In trying to understand the origin of this discrepancy, we realized that the O/H ratios of at least some of the H ii regions of the sample of Díaz & Zamora (2022) were lower than those determined by us for the same objects and clearly incorrect [^4] For example, in their supplementary tables, Díaz & Zamora (2022) give a 12+log(O/H) = 7.92 for their object number 53, which corresponds to the slit position of the Orion Nebula observed by Esteban et al. (2004), for which we obtain 8.50 (our object #422). Another examples are the H ii regions of the Magellanic Clouds observed by Toribio San Cipriano et al. (2017) or Domínguez-Guzmán et al. (2022). For those objects, Díaz & Zamora (2022) find values of 12+log(O/H) between 7.10 and 8.52 (their objects number 232 to 239), while we find a narrower –and more reasonable– range of values –between 8.02 and 8.43– for them., producing spurious higher S/O ratios. It is beyond the scope of this paper to investigate the origin and quantify the errors in the abundances calculated in Díaz & Zamora (2022), but we consider that part of their conclusions, at least those based on the chemical behaviour of the H ii regions –apparently no the SFGs– should not be taken into account until the authors provide the correct abundances. Unfortunately, the results obtained by Díaz & Zamora (2022) has led to studies exploring their origin and implications as, for example, Goswami et al. (2024), who proposed that such high S/O values might be explained by the contribution of the ejecta of Pair Instability Supernovae (PISN) produced by very massive stars ($\geq$ 130 M) in combination of a bi-modal top-heavy IMF and an initial strong burst of star formation. Ultimately, it is not necessary to invoke such exotic scenarios to explain the correct abundance patterns.
@@
In Fig. 11 we include the curves defined by the CEM of Medina-Amayo (2023) and Amayo & Carigi (in preparation) and a least-squares fit to all the data (SFGs+H ii regions) considering their errors (their parameters are listed in Table 2). The slopes of the linear fits are quite different and significantly higher for the log(S/O) vs. 12+log(S/H) distribution. The sample Pearson correlation coefficient of the log(S/O) vs. 12+log(O/H) fit is almost 0 and its p-value indicates that the distribution is indistinguishable of the null hypothesis. Therefore, we can conclude that the log(S/O) vs. 12+log(O/H) distribution of all the data is basically constant. This is not consistent with the behavior predicted by CEM, where we would expect an increase of log(S/O) of the order of +0.50 dex in the range of metallicities covered by the objects. On the other hand, the fit to log(S/O) vs. 12+log(S/H) for SFGs+H ii regions shows a clearer trend with a more positive slope, a moderate sample Pearson correlation coefficient and extremely low $p$-values (see Fig. 11 and Table 2). As we can see in Table 2, the same higher and positive slope of the fit when considering 12+log(S/H) instead of 12+log(O/H) remains when separating H ii regions and SFGs. Díaz & Zamora (2022) also find a greater dependence of the S/O ratio on S/H than on O/H in the case of SFGs, objects that apparently have correct O abundances in their paper. Fig. 11 indicates that comparison with the CEM is far more consistent in this case. The model curve fits remarkably well to the distribution of the different bins. In particular, the bins corresponding to 12+log(S/H) $\lesssim$ 6.0 show a flat trend, consistent with the behavior and the average log(S/O) given by the model. At 12+log(S/H) $\gtrsim$ 6.0, a continuous increase in S/O occurs, so the lockstep evolution of both elements is broken. Nucleosynthesis models for SNe Ia predict that these objects may produce a significant fraction of the S and Ar content in the solar neighbourhood (e.g. Iwamoto et al., 1999; Johnson, 2019; Kobayashi et al., 2020). The Chandrasekhar-mass explosion models for SNe Ia by Kobayashi et al. (2020) predict that about 29% of the total solar S originates from this process, although this fraction might be even higher if the contribution of sub-Chandrasekhar-mass SNe Ia is increased in their models. The evolution of the contribution to S enrichment of different types of stars predicted by the CEM of Medina-Amayo (2023) and Amayo & Carigi (in preparation) shown in Fig. 16, indicates that the increase in the S/O ratio with metallicity is clearly due to SNe Ia.
@@
Finally, although the linear fit to log(S/O) vs. 12+log(O/H) presents a small –but rather uncertain– positive slope that could be consistent with the hypothesis on O depletion put forward in Sect. 5.1 to explain the behaviour of the Ne/O ratio, it can not be verified in the case of S/O. Firstly because –as discussed in the previous paragraph– the production of S by SNe Ia also produces an increase in S/O with O/H. Secondly, the fraction of S depleted onto dust grains in the ISM is a controversial issue because the spectral features used to study it are often highly saturated (e.g Jenkins, 2009). In fact, it is often referred to as “the missing sulfur problem” and has raised a variety of questions over the past two decades. (e.g Keller et al., 2002; Slavicinska et al., 2024).

### v2

@@ 4.3 Argon
In Fig. 5, we plot log(Ar2+/O2+) (top) and log((Ar2++Ar3+)/O2+) (bottom) as a function of O2+/O, for H ii regions (blue circles) and SFGs (black squares) of the DESIRED-E sample. The fraction of objects with Ar2+ and Ar3+ abundances is 33.4% in SFGs, which is higher than in H ii regions (12.5%). This is because SFGs tend to have higher O2+/O ratios than H ii regions in galaxies (this is illustrated in Fig. 15). In Fig. 5, the coloured curved lines represent the behaviour of the different ICF(Ar) schemes considered: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colours corresponding to the three metallicity ranges) and Pérez-Montero et al. (2007, green dashed line). The horizontal black dashed line shows the value of the solar log(Ar/O) ratio –along as its uncertainty represented by the grey band– recommended by Asplund et al. (2021). The top panel of Fig. 5 shows that the curves of the different ICF(Ar) schemes reproduce fairly well the behaviour of the observational points except the one by Amayo et al. (2021), that does not reproduce the position of SFGs with O2+/O $\geq$ 0.9. In the case of the bottom panel of Fig. 5, we can see that all the curves reproduce quite well the distribution of the observational points. This is because the objects showing [Ar iv] lines in their spectra have necessarily a high ionisation degree and the contribution of unseen Ar+ should be very small. Moreover, the bottom right panel of Fig. 14 shows that the Ar abundance obtained with the three ICF(Ar) for the same object differs less than  0.04 dex, except for those with O2+/O $\geq$ 0.95 where the differences can be as large as 0.3 dex for those objects that lack [Ar iv] lines.
@@ 5.1 Ne/O
Summarizing the results analysed in Sec. 4.1 and Figs. 2 and 3, the ICF(Ne) by Izotov et al. (2006) shows a log(Ne/O) vs. O2+/O ratio relation flatter than the other ICF(Ne) schemes as well as the lowest standard deviation and mean and median values of log(Ne/O) (--0.67 $\pm$ 0.12 and --0.66, respectively) more consistent with the solar ratio of --0.63 $\pm$ 0.06 (Asplund et al., 2021). For these reasons we will use the Ne abundances obtained with the ICF(Ne) by Izotov et al. (2006) for the discussion we will carry out in the present section. However, even if we make this selection, determining total Ne abundances using ICFs has serious limitations. As we pointed out in Sec. 4.1, an important fraction of the objects at values of O2+/O $<$ 0.5, show very low log(Ne/O) values (see Fig. 3). In Fig. 2 it is evident that the position of those points cannot be correctly reproduced by any of the ICF(Ne) schemes considered. The strong increased scatter of the Ne2+/O2+ ratio at lower O2+/O is not something new, it was formerly reported by Kennicutt et al. (2003). Croxall et al. (2016), in their study of H ii regions in the spiral galaxy NGC 5457 (M101), confirm a similar scatter, claiming for the existence of a large population of H ii regions with significant offset to low values of log(Ne/O).
@@
Nucleosynthesis models seem to agree in predicting a basically lockstep evolution of O and Ne (e.g. Iwamoto et al., 1999; Johnson, 2019; Kobayashi et al., 2020). In Fig. 8 we also include a curve (green line) representing the evolution of Ne and O abundances of the ISM predicted by the chemical evolution model (CEM) of the Milky Way obtained by Medina-Amayo (2023). This CEM was developed for the solar neighborhood of the Galactic disc (7--9 kpc from the Galactic centre) within a two-infall gas formation scenario, similar to that of Spitoni et al. (2021). During the first 4 Gyr, the thick disc formed from a short accretion of primordial gas. Then, the thin disc began forming from a prolonged second accretion (the loop in the curve), which has continued to the present. The star formation rate and initial mass function (IMF) are based on the formulations of Kennicutt (1998) and Kroupa (2002), respectively. The model uses stellar yields for low-and intermediate-mass stars by Ventura’s group (e.g. Ventura et al., 2022, and references therein), for SNe Ia by Leung & Nomoto (2018), and for massive stars by Nomoto et al. (2013). An upper mass limit of the IMF at 37 M was inferred to match the present-day O abundance of H ii regions obtained from CELs at $\sim$8 kpc (Méndez-Delgado et al., 2022a). In Fig. 16 we show the cumulative fraction of 16O, 20Ne, 32S, 36Ar –the most abundant stable isotopes of each element, that represent between 85 to 95% of the total amount of them– produced by different types of stars as a function of 12+log(O/H) predicted by the CEM. We distinguish the fractions produced by low- and intermediate-mass stars in the asymptotic giant branch phase (AGBs, cyan line) or type Ia supernovae (SNe Ia, dark red line) and massive stars in core collapse supernovae (CCSNe, magenta line). It is important to remember that the CEM used is built to reproduce the abundance patterns observed in the Milky Way. Objects in other galaxies with a different star-formation history may be located in somewhat different places in the abundance ratios diagrams. For example, the infall of unprocessed gas into a star-forming region in a galaxy may decrease the metallicity while keeping the abundance ratio of two alpha-elements constant.
@@
Figure 8: log(Ne/O) as a function of 12+log(Ne/H) for SFGs of the DESIRED-E sample, which Ne/O ratios have been calculated using the ICF(Ne) scheme by Izotov et al. (2006). The red continuous line represents a linear fit to the data. The green continuous line shows the time evolution of Ne and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023). The dotted red line represents the mean value of log(Ne/O). The black dashed lines and the grey bands show the solar 12+log(Ne/H) and log(Ne/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
@@
As we pointed out above, in Fig. 7, the behaviour of H ii regions is very different to that shown by SFGs. In the log(Ne/O) vs. 12+log(O/H) diagram, the observational points classified as H ii regions are distributed in a quite defined range of metallicity but with a much larger dispersion of log(Ne/O). We can distinguish that although most of the objects are concentrated at log(Ne/O) values close to the solar one, there seems to be a secondary concentration at 12+log(O/H) $\sim$ 8.4 and log(Ne/O) between --1.0 and --0.7. In fact, in Table 1 the mean log(Ne/O) of the H ii regions is 0.07–0.08 dex lower than the mean of the SFGs or the solar value. Croxall et al. (2016) and Berg et al. (2020) find a similar log(Ne/O) vs. 12+log(O/H) distribution, although this was expected given that the observational data on which those studies are based are included in the DESIRED-E sample. The log(Ne/O) vs. O2+/O diagram of the bottom row of Fig. 7 shows that the dispersion towards lower values of log(Ne/O) is clearly a function of O2+/O, indicating that the ICF(Ne) is not able to correct appropriately for the presence of Ne+ in H ii regions data points at O2+/O $\leq$ 0.75, which are the vast majority of the objects of this type in the DESIRED-E sample. Kennicutt et al. (2003) suggested that this behaviour is due to the high sensitivity of Ne ionisation models to the input stellar atmosphere fluxes, which in turn are very much dependent on the treatment of opacity and stellar winds. Considering the difficulties of obtaining consistent Ne/O values from Galactic and extragalactic H ii regions, we have chosen not to take them into account for the study of the behaviour of this ratio as a function of metallicity and in the data shown in Table 2.
@@
Figure 9: log(Ne/O) as a function of 12+log(O/H). Black squares and blue circles represent SFGs and H ii regions of the DESIRED-E sample, which Ne/O ratios have been calculated using the ICF(Ne) scheme by Izotov et al. (2006). The green continuous line shows the time evolution of Ne and O abundances of the ISM predicted by the CEM of the Milky Way of Medina-Amayo (2023). Yellow and orange stars represent the abundances obtained from quantitative spectroscopic analysis of Galactic B stars by Nieva & Przybilla (2012) and Weßmayer et al. (2022), respectively. Magenta stars represent the abundances of Galactic O stars determined by Aschenbrenner et al. (2023). The black dashed lines and the grey bands show the solar 12+log(O/H) and log(Ne/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
@@
In Fig. 9 we represent the log(Ne/O) vs. 12+log(O/H) distribution of H ii regions and SFGs of the DESIRED-E sample, the curve defined by the CEM of Medina-Amayo (2023) and abundances of O and B-type Galactic stars determined from quantitative spectroscopy by Nieva & Przybilla (2012), Weßmayer et al. (2022) and Aschenbrenner et al. (2023). The atmospheres of O and B-type stars represent the present-day chemical composition of the ISM and, in principle, should provide abundance values similar to the ionised gas associated to star-forming regions. In Fig. 9, we can see that the stellar O abundances are close to solar or oversolar –as expected because most of the stars are located in the solar neighborhood– but quite separated from the locus of the bulk of the H ii regions. A systematic difference between nebular and stellar abundances of objects located in the same star-forming regions is a well-known fact. Studies comparing the O abundances of Galactic H ii regions with that of their associated O and B-type stars (e.g. Simón-Díaz et al., 2006; Simón-Díaz & Stasińska, 2011; García-Rojas et al., 2014) find stellar abundances $\sim$0.2 dex higher than nebular ones when they are calculated using CELs. However, such difference disappears when using faint recombination lines instead of CELs for deriving the nebular O abundance, this offset is the so-called abundance discrepancy problem. It may be related to the presence of temperature fluctuations inside the H ii regions (Peimbert, 1967; García-Rojas et al., 2007; Méndez-Delgado et al., 2023b). On the other hand, Fig. 9 indicates that the O and B-type stars also show a quite high dispersion in their log(Ne/O) ratio. The mean and standard deviation of log(Ne/O) of the stars included in the figure is --0.56 $\pm$ 0.17, a fairly supersolar value. Although temperature fluctuations can alter the abundances calculated with CELs, the abundance ratios –as Ne/O, S/O or Ar/O– obtained with this type of lines remain practically unchanged, so the observed log(Ne/O) offset between stellar and nebular objects is not expected to be related to this phenomenon (e.g. Esteban et al., 2020; Méndez-Delgado et al., 2024).
@@ 5.2 S/O
As in the case of the Ne/O ratio discussed in Sect. 5.1, the ICF(S) by Izotov et al. (2006) is the one that provides a more featureless –flatter– distribution between log(S/O) vs. 12+log(O/H) and vs. O2+/O (Fig. 3), a tighter fit to the (S++S2+)/O vs. O2+/O distribution in the range of ionisation degrees in common with the others ICFs (Fig. 2) and a mean log(S/O) closer to the solar value of --1.57 $\pm$ 0.05 (Asplund et al., 2021). We will adopt the results obtained using this ICF(S) for the discussion we will develop in this section. The only drawback of the ICF(S) scheme by Izotov et al. (2006) is that its application is limited to objects with O2+/O $\geq$ 0.2. In the case of the DESIRED-E sample this has a very limited impact. While no SFGs have to be removed applying this limit, the fraction of H ii regions that are finally excluded is only 13.3%. Considering the rough relation between ionisation degree and metallicity shown in Fig. 15, this implies that the removed points correspond mostly to high metallicity objects but –as we can see in the panels of the left column of Fig. 4– that area is not significantly depopulated when we compare the results using the different ICF(S) schemes.
@@
Figure 11: log(S/O) as a function of 12+log(O/H) (top) and as a function of 12+log(S/H) (bottom). Black squares and blue circles represent SFGs and H ii regions of the DESIRED-E sample, which S/O ratios have been calculated using the ICF(S) scheme by Izotov et al. (2006). The green continuous line shows the time evolution of S and O abundances of the ISM predicted by the CEM of the Milky Way of Kobayashi et al. (2020). Yellow stars represent the abundances obtained from quantitative spectroscopic analysis of Galactic B stars by Weßmayer et al. (2022). Orange stars represent the abundances of Galactic classical Cepheids determined by da Silva et al. (2023). Orange diamonds indicate the mean log(S/O) values considering bins in 12+log(S/H). The red continuous lines represent linear fits to the data represented in each panel. The black dashed lines and the grey bands show the solar 12+log(O/H), 12+log(S/H) and log(S/O) and their associated uncertainties, respectively, from Asplund et al. (2021).
@@
The behaviour of the data points of the DESIRED-E sample shown in Figs. 10 and 11 contrasts with the results obtained by some previous studies (Dors et al., 2016; Díaz & Zamora, 2022; Brazzini et al., 2024). In particular, Díaz & Zamora (2022) find that S/O vs. O/H and vs. S/H relations are quite different for SFGs and H ii regions. Díaz & Zamora (2022) report that SFGs tend to show a S/O ratio lower than solar and that it increases as the S/H increases, but remains constant with O/H, results that are in agreement with ours. However, they find that their H ii regions show S/O ratios larger than solar and a clear tendency for lower S/O ratios as O or S abundances increase, a behaviour different to what we observe. In our case, for H ii regions, since log(S/O) remains basically constant with 12+log(O/H), it shows a clear increase as 12+log(S/H) increases (see Table 2). In trying to understand the origin of the discrepancy with the results by Díaz & Zamora (2022), we realized that the O/H ratios of at least some of the H ii regions of the sample of Díaz & Zamora (2022) were lower than those determined by us for the same objects and clearly incorrect [^5] For example, in their supplementary tables, Díaz & Zamora (2022) give a 12+log(O/H) = 7.92 for their object number 53, which corresponds to the slit position of the Orion Nebula observed by Esteban et al. (2004), for which we obtain 8.50 (our object #422). Another examples are the H ii regions of the Magellanic Clouds observed by Toribio San Cipriano et al. (2017) or Domínguez-Guzmán et al. (2022). For those objects, Díaz & Zamora (2022) find values of 12+log(O/H) between 7.10 and 8.52 (their objects number 232 to 239), while we find a narrower –and more reasonable– range of values –between 8.02 and 8.43– for them., producing spurious higher S/O ratios. It is beyond the scope of this paper to investigate the origin and quantify the errors in the abundances calculated in Díaz & Zamora (2022), but we consider that part of their conclusions, at least those based on the chemical behaviour of the H ii regions –apparently no the SFGs– should not be taken into account until the authors provide the correct abundances. Unfortunately, the results obtained by Díaz & Zamora (2022) has led to studies exploring their origin and implications as, for example, Goswami et al. (2024), who proposed that such high S/O values might be explained by the contribution of the ejecta of Pair Instability Supernovae (PISN) produced by very massive stars ($\geq$ 130 M) in combination of a bi-modal top-heavy IMF and an initial strong burst of star formation. Ultimately, it is not necessary to invoke such exotic scenarios to explain the correct abundance patterns.
@@
In Fig. 11 we include the curves defined by the CEM of Kobayashi et al. (2020) [^6] The model points represented have been obtained from figures 11 and 15 of Kobayashi et al (2020) assuming their same solar abundance ratios. and a least-squares fit to all the data (SFGs+H ii regions) considering their errors (their parameters are listed in Table 2). The slopes of the linear fits are quite different and significantly higher for the log(S/O) vs. 12+log(S/H) distribution. The sample Pearson correlation coefficient of the log(S/O) vs. 12+log(O/H) fit is almost 0 and its p-value indicates that the distribution is indistinguishable of the null hypothesis. Therefore, we can conclude that the log(S/O) vs. 12+log(O/H) distribution of all the data is basically constant. This is not entirely consistent with the behaviour predicted by the CEM, where we would expect an increase of log(S/O) of the order of +0.30 dex in the range of metallicities covered by the objects. On the other hand, the fit to log(S/O) vs. 12+log(S/H) for SFGs+H ii regions shows a clearer trend with a more positive slope, a moderate sample Pearson correlation coefficient and extremely low $p$-values (see Fig. 11 and Table 2). As we can see in Table 2, the same higher and positive slope of the fit when considering 12+log(S/H) instead of 12+log(O/H) remains when separating H ii regions and SFGs. Díaz & Zamora (2022) also find a greater dependence of the S/O ratio on S/H than on O/H in the case of SFGs, objects that apparently have correct O abundances in their paper. Fig. 11 indicates that comparison with the CEM seems to be more consistent in this case. The model curve fits fairly well to the distribution of the different bins. In particular, the bins corresponding to 12+log(S/H) $\lesssim$ 6.0 show a flat trend, consistent with the behaviour and the average log(S/O) given by the model. At 12+log(S/H) $\gtrsim$ 6.0, a slight but continuous increase in S/O occurs, so the lockstep evolution of both elements is not longer achieved. Nucleosynthesis models for SNe Ia predict that these objects may produce a significant fraction of the S and Ar content in the solar neighbourhood (e.g. Iwamoto et al., 1999; Johnson, 2019; Kobayashi et al., 2020). The Chandrasekhar-mass explosion models for SNe Ia by Kobayashi et al. (2020) predict that about 29% of the total solar S originates from this process, although this fraction might be even higher if the contribution of sub-Chandrasekhar-mass SNe Ia is increased in their models.
@@
Finally, although the linear fit to log(S/O) vs. 12+log(O/H) presents a small –but rather uncertain– positive slope that could be consistent with the hypothesis on O depletion put forward in Sect. 5.1 to explain the behaviour of the Ne/O ratio, it can not be verified in the case of S/O. Firstly because –as discussed in the previous paragraph– the production of S by SNe Ia also produces an increase in S/O with O/H. Secondly, the fraction of S depleted onto dust grains in the ISM is a controversial issue because the spectral features used to study it are often highly saturated (e.g Jenkins, 2009). In fact, it is often referred to as “the missing sulfur problem” and has raised a variety of questions over the past two decades. (e.g Keller et al., 2002; Slavicinska et al., 2025).